OpenStudy (confusionist):

How to multiply radicals? Wait a moment while I type out some examples...

OpenStudy (confusionist):

\[2+\sqrt{2} (\times) 2+\sqrt{2}\]\[3+\sqrt{2}(\times)2+\sqrt{5}\]

OpenStudy (confusionist):

\[3+\sqrt{2}(\times)2+\sqrt{8}\]\[\sqrt{2}+\sqrt[3]{2}(\times)\sqrt{6}+\sqrt[3]{4}\]

OpenStudy (anonymous):

For the first example, do you mean this? \[(2+\sqrt{2})*(2+\sqrt{2})\]

OpenStudy (confusionist):

Yeah, that's what I meant for all of them. Thank you. :)

OpenStudy (anonymous):

No problem :) For future reference, the LaTeX you're looking for was this: (2+sqrt{2})*(2+sqrt{2}). Before we work on the examples, can you tell me what (a+b)(a+b) is?

OpenStudy (confusionist):

Ah, I see, cool. :) Uh, I'm not entirely sure? I've taken these from shapes that I have to find the area of. I made these equations from the given questions.

OpenStudy (anonymous):

Okay, I'll just tell you then: \[(a+b)(a+b)=a(a+b)+b(a+b)=a^2+ab+ba+b^2=a^2+ 2ab + b^2\]Do you understand this so far?

OpenStudy (confusionist):

I think so...I'm supposed to keep them in radical form as well. So I don't have to solve it necessarily.

OpenStudy (anonymous):

Don't worry, they will be :) So we know \[(a+b)(a+b)=a^2+2ab+b^2 \] Our first question was \[(2+\sqrt{2})*(2+\sqrt{2})\]If we let a=2 and b=sqrt(2), then put it into the first equation we have, can you tell me what we get out of it?

OpenStudy (anonymous):

In general, when you have something like:\[ \sqrt{2}\times 2 \]The simplified form would be:\[ 2\sqrt 2 \]

OpenStudy (anonymous):

If we have something like:\[ \sqrt{2}\times \sqrt{6} \]Then we can multiply what is in the radicals:\[ \sqrt{2\times 6} =\sqrt{12} \]However, typically we want to factor our numbers into primes:\[ \sqrt{2}\times\sqrt{2}\times\sqrt{3}= 2\times \sqrt 3=2\sqrt 3 \]